### v3

@@ 3.3.1 Technical lemmas on the problem, the framework and its minimal assumption
This section formalizes a collection of lemmas about Problem ($\mathbf{P}_{\texttt{i}}$), the POf, and Assumption 1.
@@
Third, Lemma 3 proves that the union of all partition sets of $\mathbb{Y}$ associated to indices in a neighbourhood of any $x\in\mathbb{X}$ forms a neighbourhood of all $y\in\mathbb{Y}(x)$. We use it, in our proof of Theorem 1, to map neighbourhoods of solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$) to neighbourhoods of solutions to Problem ($\mathbf{P}_{\texttt{i}}$).
@@
Then, Lemma 4 highlights several key properties resulting from the uniform continuity of $\gamma$ on each partition set of $X$ and from the finite number of partition sets of $X$ near any $x\in\mathrm{cl}(X)$, required by Assumption 1. We prove that these two properties imply that $\mathcal{ACC}(\gamma;x)$ has finitely many elements, which all belong to $\mathbb{Y}(x)$, and that any sequence in $X$ converging to $x$ has its accumulation values through $\gamma$ included in $\mathcal{ACC}(\gamma;x)$. In our proof of Theorem 1, this lemma allows us to handle the discontinuities of $\gamma$ at any solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{r}}$) identified by Algorithm 1 as an accumulation value of the sequence $(x^{k})_{k\in\mathbb{N}}$ it generates.
@@
Finally, Lemma 6 is a key result of our analysis to highlight which element of $\mathcal{ACC}(\gamma;x^{*})$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$) in our proof of Theorem 1.c).
@@ 3.3.2 Proof of Theorem 1
We now prove Theorem 1. The first two parts consist in a direct verification that given a (global or local) solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), the element $y^{*}\triangleq\gamma(x^{*})$ satisfies the definition of a usual (global or local) solution to Problem ($\mathbf{P}_{\texttt{i}}$). This is simple in the global case, while the local case uses Lemma 3 to construct the neighbourhood of $y^{*}$ involved in the definition. Then, the last part where $x^{*}$ is a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$) works similarly to the (usual) local case, except that it involves one more step to define $y^{*}$. We use Lemma 6 to carefully select $y^{*}\in\mathcal{ACC}(\gamma;x^{*})$, and then we check that $y^{*}$ satisfies the definition of a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$).
@@
Let $x^{*}$ be a global solution to Reformulation ($\mathbf{P}_{\texttt{r}}$) (so $x^{*}\in X$ so $\gamma(x^{*})\in\Omega$ exists). For all $y\in\Omega$, it holds that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$. Thus, $\gamma(x^{*})$ is a global solution to Problem ($\mathbf{P}_{\texttt{i}}$), which proves Theorem 1.a).
Let $x^{*}\in X$ (so $\gamma(x^{*})\in\Omega$ exists) be a local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$). Then there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\Phi(x)\geq\Phi(x^{*})$ for all $x\in\mathcal{X}^{*}$, and by Lemma 3 the set $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $\gamma(x^{*})$, which moreover satisfies $\chi(y)\in\mathcal{X}^{*}$ for all $y\in\mathbb{Y}(\mathcal{X}^{*})$. Thus, for all $y\in\mathbb{Y}(\mathcal{X}^{*})\cap\Omega$, it follows that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$. Thus, $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$), which proves Theorem 1.b).
Let $x^{*}\in\mathrm{cl}(X)$ be a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$). Then $\mathcal{ACC}(\gamma;x^{*})$ is finite nonempty by Lemma 4, so we may select an element $y^{*}\in\operatorname*{argmin}_{y\in\mathcal{ACC}(\gamma;x^{*})}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$. First, there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\inf\mathcal{ACC}(\Phi;x^{*})=\inf\Phi(\mathcal{X}^{*})$, and $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $y^{*}$ by Lemma 3. Second, $y^{*}$ satisfies $\inf\mathcal{ACC}(\Phi;x^{*})\geq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$ by Lemma 6. Third, $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))\geq\inf\Phi(\mathcal{X}^{*})$ holds since all $y\in\mathbb{Y}(\mathcal{X}^{*})$ satisfy $\overline{\varphi_{\Omega}}(y)\geq\Phi(\chi(y))\geq\inf\Phi(\mathcal{X}^{*})$. We deduce that $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))\geq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$, and then $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))=\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$ via Lemma 1. Thus, $y^{*}$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$), which proves Theorem 1.c). ∎
@@ 3.3.3 Proof of validity of Proposition 1 under Assumption 2
This section shows that the existing convergence analysis of Algorithm 1 (from [3, Theorem 2], and recalled in Proposition 1 above) is applicable in the context of Reformulation ($\mathbf{P}_{\texttt{r}}$). First, Proposition 2 shows that all assumptions involved in Proposition 1 are met in our context.
@@ 3.3.4 Proof of Theorem 2
Finally, we prove Theorem 2. We apply Proposition 1 to gather a solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), and then we derive a solution to Problem ($\mathbf{P}_{\texttt{i}}$) via Theorem 1.
@@
Consider Framework 1 under Assumptions 1 and 2. Let $(x^{k})_{k\in\mathbb{N}}$ be generated by Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{r}}$). Proposition 1 is applicable by Proposition 2, so all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$), so $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$) by Theorem 1.c). This proves the first part of Theorem 2. If moreover $x^{*}\in X$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$, then $\Phi$ is lower semicontinuous at $x^{*}$ by Proposition 3, so $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$) by Proposition 1, so $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$) by Theorem 1.b). This completes the proof of Theorem 2. ∎

### v5

@@ 3.3.1 Technical lemmas on the problem, the framework and its minimal assumption
This section formalizes a collection of lemmas about Problem ($\mathbf{P}_{\texttt{ini}}$), the POf, and Assumption 1.
@@
Third, Lemma 3 proves that the union of all partition sets of $\mathbb{Y}$ associated to indices in a neighbourhood of any $x\in\mathbb{X}$ forms a neighbourhood of all $y\in\mathbb{Y}(x)$. We use it, in our proof of Theorem 1, to map neighbourhoods of solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$) to neighbourhoods of solutions to Problem ($\mathbf{P}_{\texttt{ini}}$).
@@
Then, Lemma 4 highlights several key properties resulting from the completeness of $\mathbb{Y}$ and the uniform continuity of $\gamma$ on each partition set of $X$ and from the finite number of partition sets of $X$ near any $x\in\mathrm{cl}(X)$, required by Assumption 1. We prove that these two properties imply that $\mathcal{ACC}(\gamma;x)$ has finitely many elements, which all belong to $\mathbb{Y}(x)$, and that any sequence in $X$ converging to $x$ has its accumulation values through $\gamma$ included in $\mathcal{ACC}(\gamma;x)$. In our proof of Theorem 1, this lemma allows us to handle the discontinuities of $\gamma$ at any solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$) identified by Algorithm 1 as an accumulation value of the sequence $(x^{k})_{k\in\mathbb{N}}$ it generates.
@@
Finally, Lemma 6 is a key result of our analysis to highlight which element of $\mathcal{ACC}(\gamma;x^{*})$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) in our proof of Theorem 1.c).
@@ 3.3.2 Proof of Theorem 1
We now prove Theorem 1. The first two parts verify that given a usual (global or local) solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), the element $y^{*}\triangleq\gamma(x^{*})$ is a usual (global or local) solution to Problem ($\mathbf{P}_{\texttt{ini}}$). The global case is direct, while the local case uses Lemma 3 to find a neighbourhood of $y^{*}$ compatible with the definition of local optimality. The case where $x^{*}$ is a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) works similarly to the usual local case, except that it involves one more step to define $y^{*}$. We use Lemma 6 to find $y^{*}\in\mathcal{ACC}(\gamma;x^{*})$ that is a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
@@
Let $x^{*}$ be a global solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) (so $x^{*}\in X$ so $\gamma(x^{*})\in\Omega$ exists). For all $y\in\Omega$, it holds that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$. Thus, $\gamma(x^{*})$ is a global solution to Problem ($\mathbf{P}_{\texttt{ini}}$), which proves Theorem 1.a).
Let $x^{*}\in X$ (so $\gamma(x^{*})\in\Omega$ exists) be a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$). Then there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\Phi(x)\geq\Phi(x^{*})$ for all $x\in\mathcal{X}^{*}$, and by Lemma 3 the set $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $\gamma(x^{*})$, which moreover satisfies $\chi(y)\in\mathcal{X}^{*}$ for all $y\in\mathbb{Y}(\mathcal{X}^{*})$. Thus, for all $y\in\mathbb{Y}(\mathcal{X}^{*})\cap\Omega$, it follows that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$. Thus, $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$), which proves Theorem 1.b).
Let $x^{*}\in\mathrm{cl}(X)$ be a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$). Then $\mathcal{ACC}(\gamma;x^{*})$ is finite nonempty by Lemma 4, so we may select an element $y^{*}\in\operatorname*{argmin}_{y\in\mathcal{ACC}(\gamma;x^{*})}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$. First, there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\inf\mathcal{ACC}(\Phi;x^{*})=\inf\Phi(\mathcal{X}^{*})$, and $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $y^{*}$ by Lemma 3. Second, $y^{*}$ satisfies $\inf\mathcal{ACC}(\Phi;x^{*})\geq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$ by Lemma 6. Third, $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))\geq\inf\Phi(\mathcal{X}^{*})$ holds since all $y\in\mathbb{Y}(\mathcal{X}^{*})$ satisfy $\overline{\varphi_{\Omega}}(y)\geq\Phi(\chi(y))\geq\inf\Phi(\mathcal{X}^{*})$. We deduce that $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))\geq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$, and then $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))=\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$ via Lemma 1. Thus, $y^{*}$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$), which proves Theorem 1.c). ∎
@@ 3.3.3 Proof of validity of Proposition 1 under Assumption 2
This section shows that the existing convergence analysis of Algorithm 1 (from [3, Theorem 2], and recalled in Proposition 1 above) is applicable in the context of Reformulation ($\mathbf{P}_{\texttt{ref}}$). First, Proposition 2 shows that all assumptions involved in Proposition 1 are met in our context.
@@ 3.3.4 Proof of Theorem 2
Finally, we prove Theorem 2. We apply Proposition 1 to gather a solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), and then we derive a solution to Problem ($\mathbf{P}_{\texttt{ini}}$) via Theorem 1.
@@
Consider Framework 1 under Assumptions 1 and 2. Let $(x^{k})_{k\in\mathbb{N}}$ be generated by Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$). Proposition 1 is applicable by Proposition 2, so all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$), so $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) by Theorem 1.c). This proves the first part of Theorem 2. If moreover $x^{*}\in X$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$, then $\Phi$ is lower semicontinuous at $x^{*}$ by Proposition 3, so $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) by Proposition 1, so $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) by Theorem 1.b). This completes the proof of Theorem 2. ∎
4 Application of the POf to an infinite-dimensional problem
Section 1.2 mentions that the design of the POf allows addressing infinite-dimensional problems. This section proposes a simple infinite-dimensional example for illustration. Precisely we investigate a discontinuous version of a classical optimal control problem that challenges standard theoretical and numerical techniques found in the literature. In contrast, the POf allows reformulating this problem so that it may be solved analytically. This section aims for a concise analytical illustration only, using an example from the PhD thesis [11, Chapter 7.3.1]. We refer to [11, Chapter 7] and our technical report [10] for numerical experiments applying the POf to optimal control problems and using the DFPOm to solve them.
We study a discontinuous version of the textbook double integrator problem [32, 36]. This problem considers a point moving along a line over the time interval $[0,5]$, with position $p(t)$, velocity $v(t)$, and (controlled) acceleration $u(t)$ for all $t\in[0,5]$. The problem seeks to minimize the ceiling $\left\lceil{\cdot}\right\rceil$ of $p(5)$ plus the integral of $\frac{1}{2}u^{2}$. The initial state $(p,v)(0)\triangleq(0,0)$ and the final velocity $v(5)\triangleq 0$ are fixed. Let us denote also by $\mathcal{A}\mathcal{C}$ the set of absolutely continuous functions from $[0,5]$ to $\mathbb{R}$, endowed with the standard norm $\lVert{\cdot}\rVert_{\mathcal{A}\mathcal{C}}$ of the Sobolev space $W^{1,1}([0,5])\cong\mathcal{A}\mathcal{C}$, and by $\mathcal{L}^{\infty}$ the set of essentially bounded functions from $[0,5]$ to $\mathbb{R}$, endowed with the standard norm $\lVert{\cdot}\rVert_{\mathcal{L}^{\infty}}$. The problem is
$\begin{array}[]{cl}\underset{\begin{array}[]{c}(p,v,u)\in\mathcal{AC}\times\mathcal{AC}\times\mathcal{L}^{\infty}\end{array}}{\operatorname*{minimize}}&\begin{array}[]{l}\left\lceil{p(5)}\right\rceil+\displaystyle\int_{0}^{5}\dfrac{u(t)^{2}}{2}\penalty 10000\ dt\end{array}\\[4.30554pt] \mbox{subject to}&\begin{array}[t]{ll}\dot{p}(t)=v(t),&\mbox{a.e.}\penalty 10000\ t\in[0,5],\\ \dot{v}(t)=u(t),&\mbox{a.e.}\penalty 10000\ t\in[0,5],\\ (p,v)(0)=(0,0),\\ v(5)=0.\end{array}\end{array}$ ($\mathbf{P}^{\texttt{DI}}_{\texttt{ini}}$)
If Problem ($\mathbf{P}^{\texttt{DI}}_{\texttt{ini}}$) were written without the ceiling $\lceil{p(5)}\rceil$, then it would fit a usual linear-quadratic structure that could be solved using the acknowledged Pontryagin’s maximum principle (Pmp) [27]. Yet, due to the discontinuities introduced by $\lceil{p(5)}\rceil$, neither the Pmp nor its adaptations [8, 14] are applicable. Other usual techniques such as direct and indirect numerical methods [28] also fail on this problem. We refer to our technical report [10] for a more detailed discussion on this fact.
Nevertheless, as we now show, the POf allows solving Problem ($\mathbf{P}^{\texttt{DI}}_{\texttt{ini}}$) analytically. To fit the POf Framework 1, we define $\mathbb{Y}\triangleq\mathcal{AC}\times\mathcal{AC}\times\mathcal{L}^{\infty}$, and the norm $\lVert{(p,v,u)}\rVert_{\mathbb{Y}}\triangleq\lVert{p}\rVert_{\mathcal{A}\mathcal{C}}+\lVert{v}\rVert_{\mathcal{A}\mathcal{C}}+\lVert{u}\rVert_{\mathcal{L}^{\infty}}$ for all $(p,v,u)\in\mathbb{Y}$. Then we define the cost functional $\varphi:y\triangleq(p,v,u)\in\mathbb{Y}\mapsto\left\lceil{p(5)}\right\rceil+\frac{1}{2}\int_{0}^{5}u(t)^{2}\penalty 10000\ dt$. To partition $\mathbb{Y}$, we define $\mathbb{X}\triangleq\mathbb{R}$ and for all $x\in\mathbb{X}$, we set $\mathbb{Y}(x)\triangleq\{(p,v,u)\in\mathbb{Y}:p(5)=x\}$. It follows that $\chi$ is defined as $\chi(p,v,u)\triangleq p(5)$ for all $(p,v,u)\in\mathbb{Y}$. Then we apply the POf as described in Section 3.1. For all $x\in\mathbb{X}$, the subproblem of minimizing $\varphi$ over $\mathbb{Y}(x)$ reads as
$\begin{array}[]{cl}\underset{\begin{array}[]{c}(p,v,u)\in\mathcal{AC}\times\mathcal{AC}\times\mathcal{L}^{\infty}\end{array}}{\operatorname*{minimize}}&\begin{array}[]{l}\left\lceil{p(5)}\right\rceil+\displaystyle\int_{0}^{5}\dfrac{u(t)^{2}}{2}\penalty 10000\ dt\end{array}\\[4.30554pt] \mbox{subject to}&\begin{array}[t]{ll}\dot{p}(t)=v(t),&\mbox{a.e.}\penalty 10000\ t\in[0,5],\\ \dot{v}(t)=u(t),&\mbox{a.e.}\penalty 10000\ t\in[0,5],\\ (p,v)(0)=(0,0),\\ (p,v)(5)=(x,0).\end{array}\end{array}$ ($\mathbf{P}^{\texttt{DI}}_{\mathrm{sub}}(x)$)
For all $x\in\mathbb{X}$, Subproblem ($\mathbf{P}^{\texttt{DI}}_{\mathrm{sub}}(x)$) has a classical linear-quadratic structure, since $\lceil{p(5)}\rceil=\lceil{x}\rceil$ is constant. Thus, Subproblem ($\mathbf{P}^{\texttt{DI}}_{\mathrm{sub}}(x)$) admits a solution and the Pmp is applicable. This latter provides the necessary optimality condition that, if $(p^{*},v^{*},u^{*})\in\mathcal{AC}\times\mathcal{AC}\times\mathcal{L}^{\infty}$ is a solution to Subproblem ($\mathbf{P}^{\texttt{DI}}_{\mathrm{sub}}(x)$), then there exists a nontrivial triplet $(q_{0}^{*},q_{p}^{*},q_{v}^{*})\in\{-1,0\}\times\mathcal{AC}\times\mathcal{AC}$ such that
[…]
Then, for all $x\in\mathbb{X}$, elementary calculations ensure that System (Pmp) has a unique solution given by
$(p^{*}_{x},v^{*}_{x},u^{*}_{x})(t)=\dfrac{x}{125}\left(t^{2}(15-2t),t(30-6t),30-12t)\right)\quad\mbox{a.e.}\penalty 10000\ t\in[0,5],$
and thus, Subproblem ($\mathbf{P}^{\texttt{DI}}_{\mathrm{sub}}(x)$) has a unique global solution given by
$\gamma(x)\triangleq(p^{*}_{x},v^{*}_{x},u^{*}_{x}),\quad\text{with}\quad\varphi\left(\gamma(x)\right)=\left\lceil{x}\right\rceil+\dfrac{12}{125}x^{2}.$
Hence, in this context, Reformulation ($\mathbf{P}_{\texttt{ref}}$) reads as
$\underset{\begin{array}[]{c}x\in\mathbb{X}\end{array}}{\operatorname*{minimize}}\quad\Phi(x),\quad\text{where}\quad\Phi\triangleq\varphi\circ\gamma.$ ($\mathbf{P}^{\texttt{DI}}_{\texttt{ref}}$)
Now that Problem ($\mathbf{P}^{\texttt{DI}}_{\texttt{ini}}$) fits into the POf Framework 1, observe that Assumption 1 holds. Indeed, $\chi$ is continuous, and $\gamma$ is uniformly continuous so Assumption 1 is nonstringent with respect to the partition of $\mathbb{X}$. A natural admissible partition of $\mathbb{X}$ is given by $\mathbb{X}=\sqcup_{n\in\mathbb{Z}}]n,n+1]$, since the ceiling $\lceil{x}\rceil$ involved in $\varphi(\gamma(x))$ is constant over each of these intervals. Consequently, Theorem 1 is applicable. By Theorem 1.a), any global solution $x^{*}$ to Reformulation ($\mathbf{P}^{\texttt{DI}}_{\texttt{ref}}$) leads to $\gamma(x^{*})$ as a global solution to Problem ($\mathbf{P}^{\texttt{DI}}_{\texttt{ini}}$). Moreover, the minimizer of $\Phi$ is tractable as $x^{*}=-5$. Thus, by Theorem 1.a),
[…]
This proves that the discontinuous Problem ($\mathbf{P}^{\texttt{DI}}_{\texttt{ini}}$), which challenges standard techniques from the literature on optimal control theory, is solved analytically using the POf. Remark that a solving a naive alteration of Problem ($\mathbf{P}^{\texttt{DI}}_{\texttt{ini}}$) that consists in replacing $\lceil{p(5)}\rceil$ by $p(5)$ in the objective function fails to recover a global solution to the true problem. Indeed, this altered problem is tractable using the Pmp, but its solution equals $\gamma(x^{\sharp})$ with $x^{\sharp}\triangleq\frac{-125}{24}\approx-5.20\neq x^{*}$. This example illustrates how the POf may be used to tackle some infinite-dimensional problems that admit difficulties depending on a few variables. Moreover, we stress that we need neither a discretization scheme to make the dimension finite nor a DFO algorithm to solve Reformulation ($\mathbf{P}^{\texttt{DI}}_{\texttt{ref}}$), since we obtain its solution analytically. Thus, this examples relies only on the POf and not on the DFPOm, so we do not need to check Assumption 2.